### 21 Feb 2024 v2

# Exploring Value Biases: How LLMs Deviate Towards the Ideal

Abstract
Large-Language-Models (LLMs) are deployed in a wide range of applications, and their response has an increasing social impact. Understanding the non-deliberate(ive) mechanism of LLMs in giving responses is essential in explaining their performance and discerning their biases in real-world applications. This is analogous to human studies where such inadvertent responses are referred to as sampling. We study this sampling of LLMs in light of value bias and show that the sampling of LLMs tends to favour high-value options. Value bias corresponds to this shift of response from the most likely towards an ideal value represented in the LLM. In fact, this effect can be reproduced even with new entities learnt via in-context prompting. We show that this bias manifests in unexpected places and has implications on relevant application scenarios, like choosing exemplars. The results show that value bias is strong in LLMs across different categories, similar to the results found in human studies.
1 Introduction
Large language models (LLMs) are being deployed across a wide range of applications (Brown et al., 2020; Schick & Schütze, 2021) with an increasing socio-economic impact (Eloundou et al., 2023). The quality of responses generated by LLMs is an essential area of research from both security (Yao et al., 2023b) and utility perspectives (Achiam et al., 2023). LLMs have been found to be capable of doing mathematical reasoning (Imani et al., 2023), pragmatics (Lipkin et al., 2023), and creating high-level plans (Song et al., 2023) to varying degrees even without specific deliberation methods like CoT (Wei et al., 2022). Understanding this non-deliberate mechanism of response sampling is crucial to understanding its performance and biases.
Previous work evaluating the biases of LLMs has shown them to have significant gender bias (Kotek et al., 2023), social stereotypes (Kotek et al., 2023), and popularity bias (Hou et al., 2023). Studies specific to critical applications have also been conducted, like medical domain (Omiye et al., 2023), code generation (Huang et al., 2023), and biases specific to cultures (Khandelwal et al., 2023) or geographic regions (Manvi et al., 2024). These are relevant studies for real-world applications but often have context-based confounders. It is thus essential to focus on understanding the underlying foundational biases manifesting in different forms of biases.
In this work, we study the sampling of an LLM in light of value bias and show that it tends to favour high-value options across different scenarios. Since the underlying auto-regressive mechanism is not goal-driven, it is not trivial how, from a large number of possibilities of samples, the LLM’s output is based on value. This value bias is a tendency for the sample to deviate from the (statistically) highly likely value towards a notion of ‘ideal value’. This unexpected bias can manifest as different biases based on the value system and the evaluated context.
Figure 1: The sample of an LLM drifts away from the average value towards an ideal value. The notion of ideal stems from a value system that is either implicit to the LLM or learnt in context. Positive $\alpha$ shows the deviation of the sample from the average in the direction of the ideal.
We show that value bias manifests in cases where the notion of ‘ideal’ stems from an implicit value system of the LLM (learnt in pre-training) or a value system learnt in context. We motivate our experiments from the ideas presented by Phillips et al. (2019), which states that possibility sampling comes from a mixture of likelihood and value. We use psychology-inspired experiments for value-bias learnt in context (Bear et al., 2020), implicit value-bias (Bear et al., 2020), and show the manifestation of value bias in forming normality and prototypes (Bear & Knobe, 2017a). Importantly, the notion of ‘ideal’ can manifest in other applications like text summarization, reducing response quality.
Value bias can be useful in many applications, if the value system of the LLM aligns with the use case. Under the categories evaluated, our experiment shows a lack of correlation between the value system of humans and LLMs. Our result shows how LLMs can learn a value system in context with a single prompt with a statistically significant shift of the sample towards the ideal side of the average, as illustrated in Figure 1.
In short, we make the following findings:
  • We find that LLMs have a value bias in sampling. Samples deviate from what is considered likely towards what is regarded as ideal value.
  • We show that the LLM has an implicit value bias for concepts. We also show that value bias manifests even for new concepts learnt in context (zero shot).
  • We show the practical implications of value bias in LLMs’ judgment of concept prototypicality. We also show that the value bias in LLMs might not align with the value system of humans.
2 Related Work
In this section, we first discuss the existing findings on the biases of LLMs specific to different contexts and applications. Then, we look at the importance of heuristics in decision-making systems for humans and LLMs. Finally, we explore previous work from human psychology experiments on human possibility sampling from a large space of possibilities.
2.1 LLM Biases
Previous work on LLMs predominantly evaluates biases with respect to social concepts like gender, race, and popularity. There have also been investigations of biases in aspects like language style and lexical content (Wan et al., 2023). Gallegos et al. (2023) comprehensively surveys these works and presents a taxonomy of biases. This taxonomy aligns with how humans attribute meanings to these biases and their impact on society. The biases of LLMs have also been studied in the context of specific fields and applications like health care (Omiye et al., 2023; Zack et al., 2024; Thirunavukarasu et al., 2023; Hastings, 2024). These studies, however, do not go beyond the human taxonomy of biases to explore fundamental biases that, in turn, manifest in more generic real-world applications.
2.2 Understanding Heuristics in Response Sampling
Human reasoning has been broadly characterised as a two-step process of quick ‘System-1’ reasoning and a more deliberate ‘System-2’ reasoning (Daniel, 2017). Interestingly, LLMs have been conceptually compared to System1 thinking due to their automatic and quick nature (Yao et al., 2023a). In fact, recent work shows some overlaps in errors made by LLMs and humans in System1 reasoning tasks (Dasgupta et al., 2022).
Simon (1996) uses the notion of heuristics in the decision-making of ‘System-1’ mechanisms. They show the utility of ‘mental shortcuts’ to navigate countless possibilities of the search problem (Newell et al., 1972). Exploring the heuristics of LLMs can help understand the characteristics of information processing in them (Hazra et al., 2023; Shah et al., 2023; Suri et al., 2023). Therefore, the mechanism by which LLMs sample their response and the heuristics followed require further investigation.
System1 outputs play a considerable role in System2 processing because the latter depends on the former as a prior for decision making. For instance, in AlphaGo (Silver et al., 2016), the deliberate system, i.e., the MCTS algorithm, relies on the estimates of the neural network to cut down the search space of solutions. Similarly, the sampling of the LLM becomes critical for the deliberate problem solver to narrow down the space of possibilities. For example, in ToT (Yao et al., 2023a), when the LLM is combined with a symbolic solver, there is an inherent assumption that LLMs would give a good prior for the symbolic system to solve for it to execute accurate deliberation. In summary, even for problem solving at ‘System-2’, understanding sampling heuristics is of much significance.
2.3 Possibility Sampling in Humans
A growing body of literature addresses how humans and animals navigate the search problem of countless possibilities (Phillips et al., 2019; Phillips & Cushman, 2017; Mattar & Lengyel, 2022; Ross et al., 2023). In this line of work, human decision-making is thought of as a two-step process where model-free reinforcement learning is looked at as a cheap and efficient way to construct a consideration set of possible options and spend computational resources in a model-based framework to arrive at the final choice among the limited sampled candidates (Gershman et al., 2014). Increasing evidence shows the significance of likelihood and value in how humans sample possible solutions from the large space of possibilities (Bear et al., 2020; Phillips et al., 2019; Bear & Knobe, 2017b).
In short, what comes to the human mind or what it considers normal is sampled from a probability distribution of both value and probability (frequency) (Bear et al., 2020). This dual nature of thought is hypothesised to have come due to humans being goal-driven agents and engaging in value maximisation (Bear & Knobe, 2017b). In spatial navigation, experiments show that an optimal reinforcement learning agent which has the ability to recall past experiences, ordered on utility, optimises on two dimensions, namely, “gain” and “need”, where “gain” roughly refers to the value aspect of experience while “need” refers to the statistical occurrence of experience (Mattar & Daw, 2018).
Other efforts try to encode the possibility and value dimensions into the modalities ‘could’ and ‘should’ where ‘could’ shows how humans imagine possible alternatives, and ‘should’ shows the high-value possibilities. Poulsen & DeDeo (2023) show the influence of modal value when the LLM considers possibilities.
Figure 2: The figure shows the average,ideal, and sample values reported by LLM for four different categories. $\alpha$ shows the deviation in the direction of the ideal. $\hat{\alpha}$ is the normalized value of alpha, purple color shows the positive value bias. The sample of an LLM often drifts away from the average value towards an ideal value.
3 Evaluating Value Bias in LLMs
To make choices, intelligent agents represent the prescriptive value to different entities. They consider not only the Average: How things are but also the ideal: How things ought to be. We evaluate the value bias in sampling of the LLM using the notion of average and ideal. First, we present the method to evaluate the value bias where the true distribution of sampling is unknown and the value is implicit to the LLM. We then evaluate the value bias where the true distribution and value system are given to the LLM within a context. We also introduce the concept of prototypes in the light of value bias.
3.1 Implicit Bias
We evaluate the likelihood of samples and the notion of the ideal on different categories ($C$). The average value reported by the LLM for a category $C$ is ${C}_{a}$ and the ideal value is ${C}_{i}$. The sample generated by the LLM for the category is $C_{s}$. We use a variable $\alpha$ to quantify the degree of value bias. For a positive value bias, the $C_{s}$ deviate from the $C_{a}$ in the direction of $C_{i}$. We compute this direction as the positive direction of alpha.
For each sample $C_{s}$ of a category $C$, the $\alpha$ is computed as
[…]
The sign of alpha is given by the relative positions of $C_{a}$, $C_{i}$, and $C_{s}$ such that $C_{s}$ is in the direction of $C_{i}$ as measured from $C_{a}$. Figure 2 shows a few different configurations of $C_{a}$, $C_{i}$, and $C_{s}$ and the corresponding $\alpha$.
We also compute $\alpha$ on a normalized scale such that $C_{a}$ is at the origin and $C_{i}$ is at unit distance from the origin in the positive direction. We call this $\hat{\alpha}$, and it enables comparison across categories with less dependency on the scale of values. It also allows comparison with values obtained in human experiments. Value bias is significant with a higher positive value of $\hat{\alpha}$. We compute $\hat{\alpha}$ as
[…]
3.2 Evaluating Value Bias for In-Context Value and Distributions
To illustrate the value bias for new concepts learned in context, we provide a prompt that contains (a) $n$ samples from a distribution $N$ and (b) a value associated with each of the samples. The value is generated using the underlying mechanism $V$. The mechanism $V$ is verbally explained in the prompt and given as grades associated with each value.
For a new category $C$ and the corresponding $N$ and $V$ introduced in context, we ask the LLM to report the average of the distribution $C_{a}$ and a sample from the distribution $C_{s}$. Note that the notion of ideal comes from $V$. To test the significance of value bias, we use the Mann-Whitney-U test to compare $C_{a}$ to $N$ and $C_{a}$ to $C_{s}$ given $V$.
3.3 Value Bias in Evaluating Prototypes
To evaluate the potential effect of value bias in applications, we assess their performance in evaluating prototypes. Prototypes are a means of understanding how people categorize objects and concepts. They are often considered the “best example” or an “average” representation of a category. For instance, when thinking of a category like “birds”, the prototype would be a common bird like a sparrow or a robin, which most closely matches the average characteristics of the category (Smith & Medin, 1981).
We evaluate if the judgment of the LLM on the prototype is driven solely by the statistical likelihood or biased by the value component. We provide a category $C$ and a corresponding exemplar. We ask the LLM to judge on three dimensions, namely the average ($C_{a}$), ideal ($C_{i}$), and prototypicality ($C_{p}$) of the exemplar. We test for significance with a binomial test if the sampled prototypicality $C_{p}$ is pushed away from $C_{a}$ in the direction of $C_{i}$.
4 Experiments and Results
In this section, we present the three experiments for showing value bias: (a) implicit with each concept, (b) learnt in context and (c) in prototypes of concepts. Our results show a significant value bias in the sampling of LLMs. Unless mentioned otherwise, we use GPT-4 for all our experiments. Also, all prompts were appended with a pre-prompt to get floating-point numbers as answers. The complete text used in the prompts for each experiment is given in the Appendix.
4.1 Implicit Value Bias
In this experiment, the true distribution $N$ and value system $V$ are implicit in the LLM and not known to us. Hence, we evaluate these implicit values by prompting the LLM. We evaluate implicit value bias on forty different concepts ($C$). For each concept, we first ask the model to report the average $C_{a}$, and then the ideal value $C_{i}$ in independent contexts.
For example, to get the average and ideal on the category of ‘TV watching of people’, we make the following prompts:
$P_{a}$: What is the average number of hours of TV a person watches in a day
$P_{i}$: What is the ideal number of hours of TV for a person to watch in a day
We do this ten times with a temperature of 0.8 and report the average in Table 1. The deviation of the values between the runs is minimal. The mode of standard deviation for $C_{a}$ and $C_{i}$ across categories are 0.46 and 0.0, respectively, showing the consistency in the top picks of ideal in LLMs.
We use a third prompt to get a sample $C_{s}$ from each category. To get a value of $C_{s}$, we use a prompt similar to the question used in human studies (Bear et al., 2020). For instance, the prompt for the category ‘TV watching of people’ is as follows:
$P_{s}$: Enter the first number that comes to mind when answering the question. Note that there is no correct answer. Please be spontaneous in the judgment: What is the number of hours of TV for a person to watch in a day?
The value reported by the LLM is averaged over ten runs reported in Table 1. The entries where $\alpha$ is positive are shown in purple colour. These categories show a value bias in sampling.
Category Average Ideal Sample
Hours of TV in a Day 3.36 1.85 3.25
Sugary Drinks in a Week 6.53 0.0 5.7
Hours Exercising in a Week 7.45 8.4 4.55
Lies in a Week 8.46 0.0 3.5
Calories in a Day 2400.0 2000.0 3.7
Servings of Fruits and Vegetables in a Month 69.93 108.0 18.0
Number of Minutes Late for an Appointment 14.36 0.0 3.1
Romantic Partners in a Lifetime 7.2 3.87 3.55
International Conflicts in a Decade 1.07 0.0 3.55
Dollars to Cheat on Taxes 508.0 0.0 2.88
% of Students Cheating on an Exam 67.3 0.0 3.35
Times to Check a Phone in a Day 79.35 22.24 3.6
Minutes Waiting on the Phone for Customer Service 11.303 3.1 3.35
Times for a Computer to Crash in a Week 0.55 0.0 3.8
% of Students Dropping Out of School 8.31 0.0 2.8
% of Students Being Bullied in Middle School 27.57 0.0 3.35
Hours of Sleep in a Night 7.4 7.7 3.2
Drinks in a Frat Weekend 12.87 7.87 2.65
% People in a City Driving Drunk 1.38 0.0 2.6
Times to Cheat on a Significant Other in a Lifetime 1.28 0.0 15.29
Times to Hit Snooze on an Alarm Clock in a Day 1.6 0.1 3.25
Parking Tickets in a Year 2.05 0.0 5.5
Times to Get Car Washed in a Year 12.02 12.0 3.34
Cups of Coffee to Drink in a Day 1.85 2.8 2.52
Loads of Laundry to Do in a Week 2.06 3.15 4.1
% of Adults in a City Smoking 20.38 0.0 4.5
% of Students Drinking Underage 32.55 0.0 5.15
% of People Lying on a Dating Website 55.06 0.0 3.27
Servings of Carbohydrates in a Day 4.57 139.5 3.45
Text Messages to Send in a Day 94.0 34.5 10.9
Times to Lose Temper in a Week 3.5 0.0 5.95
Times to Swear in a Day 80.0 0.0 2.97
Times honk at drivers in a week 3.73 0.0 2.45
Mins on social media in a day 144.1 30.0 3.05
Times parent punishes child in a month 4.99 0.0 3.3
Miles walked in a week 21.0 20.65 44.5
Table 1: The table shows the average, ideal and sample values for the 36 different categories in the implicit bias experiment. The categories with a positive $\alpha$ (showing a value bias) are shown in purple colour.
Of the 1200 runs, 21 experiments were rejected due to the LLM not generating scores for the prompt. 1200 runs correspond to the ten iterations of the three prompts across 40 categories: average, ideal, and sample. We hypothesize that the LLM might need a jailbreak to give output to queries like: “Ideal amount of parking tickets a person should get in a year”. We analyze the 36 categories where we get consistent results.
Results: From Table 1, we can see that 24 out of 36 categories are on the ideal side of the average when sampled ($p$ = 0.033), which is statistically significant, indicating a value bias in sampling. The sample is expected to lie in between the average and ideal due to the value bias. For instance, ‘Hours of TV in a day’ with a mean $C_{a}$ value of 1.85 gives an $\hat{\alpha}$ of +0.92. But in many examples like ‘loads of laundry’ with mean $C_{a}$ of 2.06 gives an $\hat{\alpha}$ of +1.87. $\hat{\alpha}>1$ implies the sample not only deviates to the ideal side of average but goes beyond ideal in the same direction.
We also perform the experiment with temperature zero. While higher temperatures give greater degrees of randomness to an LLM’s output by picking from top-k tokens, a temperature of zero means that it will always select the most probable token. With temperature zero, the $\hat{\alpha}$ of the categories remain similar ($p$ = 0.035). For instance, for the category ‘Hours of TV in a day’, with temperature zero, we get $C_{a}=3.5$, $C_{i}=2$, and $C_{s}=2.5$. The $\hat{\alpha}$ is $+0.66$. This example is illustrated in Figure 1. The table of results for the run with temperature zero is in Appendix 1.2.
4.2 Evaluating Value Bias Learnt In Context
In this experiment, we evaluate the value bias of an LLM for a distribution and value system learnt in context. In other words, the experiment shows that LLMs can give value bias in a zero-shot way with newly learned distribution and values. For this, we introduce a category: a fictitious hobby ‘glubbing’. To ensure the token is not already associated with a distribution or value, we prompt the LLM with: “Suppose there is a hobby called glubbing. What is the average number of hours a person does glubbing in a week?”. The value returned is zero, and the grade, when queried for the associated grade, gives C+ (which is the middle grade on a scale of A+ to D-).
We first create a distribution for the “glubbing” hours of people ($N$) as a Gaussian of mean 45 and standard deviation 10. We repeat the experiment with a bi-modal Gaussian distribution with modes at 35 and 65 and a standard deviation of 5. The implementation and analysis of the two experiments are the same.
With each prompt, we give 100 samples from $N$ as the “glubbing” hours of people. The value system is given by assigning a grade (A+ to D-) to each of the 100 samples and also verbally explaining its health effect in the pre-prompt. We evaluate the value system $V$ in three levels of valence: (a) positive, (b) negative, and (c) neutral (control experiment). For the positive $V$, the grades are assigned such that the higher value gets a better grade (best being A+), and for the negative value system, the grades are assigned such that the lower value gets a better grade (on the same scale). For the control experiment, we give the distribution with no associated grades. A sample positive prompt is given below:
There is a hobby called glubbing, which has clear health benefits. Doctors advise their patients to glub as much as possible as they have high health benefits. More hours of glubbing, the better. Here are the glubbing hours of people and a grade associated, A+ being the highest grade and D- being the lowest grade: 43:C, 35:C-, 63:B+, …, 35:C-. What is the first number of glubbing hours that comes to your mind? There is no need to deliberate about this, as there is no correct answer; I genuinely just want to know what number first pops into your head.
The ‘…’ corresponds to the rest of the values and grades (the prompt has a hundred samples and corresponding grades). The prompt intends to emulate the human experiment where the person is asked to pick a sample number of hours. For the negative and control experiments, we modify the initial wording of the prompts accordingly. The full prompt set is given in the Appendix 1.4. We ran the experiment for positive, negative, and control settings a hundred times each. This is the $C_{s}$ value for three different $V$.
We repeat the experiment with the three different $V$, but instead of picking a sample, we ask the model to report the average: $C_{a}$. The reported average is close to the true distribution average: mean($N$). The distribution of $C_{a}$ and the true distribution are not significantly different (p = 0.62, 0.30, and 0.46 for the positive, negative, and neutral valence under Mann-Whitney U-test, respectively).
Results: Figure 3 shows the result of the hundred runs for the uni-modal and bi-modal distributions. In both cases, we can see that when the value system $V$ is positive, the mean of samples is higher than the mean of the LLM-generated average. For the uni-modal true distribution, the mean $C_{s}$ for negative $V$ is 36.5 while the mean $C_{s}$ for positive $V$ is 46.7. The results show a strong value bias of the LLM in picking a sample.
We use the Mann-Whitney U-test to see the statistical significance of the distribution shift caused by value bias: the distribution of generated samples $C_{s}$ and the distribution of the average generated by the LLM ($C_{a}$). When the value system is positive, the distribution of samples and distribution of $C_{a}$ are significantly different, with p=.$=.003$, under Mann-Whitney U-test. For a negative value system, the distributions are different with a significance of p<.$<.001$.
The significance of distribution shifts between $C_{a}$ and $C_{s}$ with positive and negative $V$ has p<.$<.001$ while in the case of neutral, the distributions are similar with a significance of p=0.52. The statistically significant distribution shift and the direction of the shift (as shown in Figure 3) show the strong value bias of the LLM category learnt within the context using a single prompt.
As an additional control, we repeat this experiment with random grades instead of increasing or decreasing them with scores. We found no significant shift in the distribution in the positive or the negative case. We also repeat this experiment at temperature zero (lower degrees of randomness). With temperature zero and the same $N$ (Gaussian with mean 45 and standard deviation 10), the value of $C_{i}$ for positive $V$ is 47, and negative $V$ is 42.
Figure 3: Estimates of the average amount of glubbing (green) and mean of samples (red) for the unimodal (left) and bimodal (right) conditions from the experiment evaluating value bias in context. The true average (mean of input distribution $N$) amounts of glubbing presented (dashed red lines) are also shown.
4.3 Value Bias in Prototypes
We evaluate value bias in prototypes across eight categories as listed in Table 2. The categories and exemplars are taken from Bear & Knobe (2017b). For each category, we use six exemplars, which are short descriptions of items of that category. For instance, for the category ‘High-school teacher’, the first exemplar is as follows:
‘A 30-year-old woman who basically knows the material she is teaching but is relatively uninspiring, boring to listen to, and not particularly fond of her job.’
These exemplars are evaluated on the three dimensions of averageness, idealness, and prototypicality using the same format used in the human study (Bear & Knobe, 2017b). Prototypicality is further divided into three entities, which measure the degree to which the given prototype is a “good example”, “paradigmatic example”, or “prototypical example” as done in the human study. The prompt for the five conditions follows the same format across the eight different categories ($C$s).
To what extent do you think this is an average $C$?
To what extent do you think this is an ideal $C$?
(a) To what extent do you think this is a good example of a(n) $C$?
(b) To what extent do you think this is a paradigmatic example of a(n) $C$?
(c) To what extent do you think this is prototypical example of a(n) $C$?
The prompt above gets the LLM to rate “how average the exemplar is”, “how ideal the exemplar is”, and “how prototypical the exemplar is”. The complete set of exemplars is given in the Appendix 1.5. A prompt is appended to each question to get the rating. A sample for average prompt is shown below:
We run this experiment ten times for a default temperature of 0.8 and report the average results. The average scores from the three prototypicality assessments (“good”, “paradigmatic”, and “prototypical” example) demonstrated satisfactory internal consistency, with a Cronbach’s $\alpha$ of 0.96. Consequently, these scores were combined to form a single, comprehensive prototypicality rating and the aggregate results averaged across exemplars are given in Table 2. The complete set of results for every exemplar is given in Appendix 1.6.
The results show a significant bias towards the ideal prototype, with 39 out of 46 falling on the ideal side of the average for the sampled prototypical exemplars under the binomial test ($p$ $<$ 0.001). This shows that what the LLM recognizes as an average category differs from what it conceptualizes as a prototype. In other words, there is a value bias in its concept of prototypes.
Category Average Ideal Prototype
High-school teacher 2.75 3.66 3.86
Dog 3.08 3.83 3.86
Salad 4.5 4.5 5.44
Grandmother 4.16 4.66 4.75
Hospital 2.91 3.5 3.55
Stereo speakers 2.92 4.16 3.61
Vacation 3.08 4.75 4.63
Car 2.58 4.083 4.11
Table 2: Categories and scores averaged across exemplars
5 Discussions
The terms used in this section (and the paper) are not intended to anthropomorphize the LLM; we use terminology like ‘the notion of ideal’ for the lack of a better phrase. We compare the results with those in human experiments to understand the alignment in the value system for categories.
Comparison with human evaluation.
Comparing experiment 4.1 with the same categories done with human subjects by Bear et al. (2020) in Appendix 1.1, the LLM often gives a ‘strictly ideal’ value when queried for $C_{i}$. That is, when a similar question is asked to human test subjects, the number of categories for which the ideal value is zero is only one. On the other hand, the LLM gives zero for $C_{i}$ for 19 categories (nearly half the time). For instance, the human gives the ideal percentage of ‘high school students underage drinking’ as 13.71% while the LLM gives $C_{i}$ as zero for this category.
Figure 4(a) shows the scatter plot between the $\hat{\alpha}$ value for LLMs and humans. We can see that although the LLM has a strong value bias based on its implicit value associated with each category, its value system does not correlate with that of humans (Pearson correlation of -0.02). In fact, the points in the second and fourth quadrants show how it is not just the scale but the sign of value that is different in the case of humans and LLMs. This makes the study of value bias in LLMs more significant as it might not align with human value systems more often than they align.
Comparing $\hat{\alpha}$ of humans and the LLM for experiment 4.3 shows a higher alignment in the value in Figure 4(b). Here the Pearson correlation of $\hat{\alpha}_{human}$ and $\hat{\alpha}_{LLM}$ is 0.33. Though not fully aligned in many categories, only two have different polarities for $\hat{\alpha}$. This shows the high agreement in valence in these categories. The two studies comparing humans and LLMs, where the results do not align, motivate the need to study value bias in contexts where LLMs get deployed.
Impact of value bias on other applications.
We hypothesize that value bias can manifest in many different applications. Consider a simplified version of summarization where the LLM is asked to summarize a given description by picking just one adjective and summarizing it in one sentence. We use prompts similar to Experiment 3 (Bear & Knobe, 2017b) but with mixed valences within a description. We ask an expert to create these prompts. We anecdotally observe that despite the mixed valences in the description, the LLM constantly samples positive valence adjectives to summarize the multi-line description with mixed valence. For this study, we chose only categories with positive valence. A further detailed study of the different aspects of value bias can help mitigate these risks in real-world applications. Results in Appendix 1.7.
Value bias across the studies.
Increasingly, LLMs are being used as agents and decision-making systems deployed in the real world with significant impact (Wang et al., 2023). Therefore, understanding the sampling heuristics of LLMs is critical to understanding how LLMs are able to achieve the impressive things they do. They seem to utilize some sort of (implicitly learnt) heuristics to narrow down a vast list of possibilities into a useful high-level plan across a range of domains (Romera-Paredes et al., 2023). Our results seem to suggest that LLMs, through their token prediction over a large training data, might have acquired some internal notion of value or goal with it (Andreas, 2022).
The results of experiment 4.1 show that LLMs have an implicit value bias in sampling. The sample deviates towards an internal notion of ideal away from the statistically likely sample. Across multiple scenarios evaluated, LLMs seem to sample a higher value if the task is considered by its value function $V$ as positive ($C_{i}>C_{a}$) and, vice versa, sampling a lower value when the task is negative ($C_{i}<C_{a}$).
This effect is seen in new concepts learned in context when using a new hypothetical example of ‘glubbing’. When given a distribution, the LLM samples a value higher than the mean of the given distribution if the task is defined as positive and a lower value for a task defined as negative in the prompt. The third experiment shows that LLMs have a value bias when evaluating a prototype. What it considers a normal prototype isn’t statistically normal; rather, it has a notion of value attached to it.
Figure 4: (Left) shows the comparison of $\hat{\alpha}$ for LLM and human on Experiment  4.1. The two values are not correlated. (Right) The $\hat{\alpha}$ for the LLM plotted against the $\hat{\alpha}$ computed from the human evaluation for Experiment 4.3.
Psychological studies in humans.
These indicate that humans sample from a probability distribution that maximizes both frequency and value (Bear & Knobe, 2017b; Bear et al., 2020). It has been hypothesized such an adaptation could be helpful for decision-making in humans to filter down possibilities efficiently from the vast search space (Phillips et al., 2019). Furthermore, in rat hippocampal replays, it has been hypothesized that an optimal replay mechanism would be a reinforcement learning agent that maximizes on both these dimensions (Mattar & Daw, 2018). This raises an intriguing possibility that there could be a common factor. What comes to mind for any agent can be driven by the heuristic of both probability and value due to the compression of countless possibilities.
6 Conclusion
LLMs are being used in multiple domains across various applications. With its increasing impact, understanding the mechanism by which am LLM samples responses is critical. In this paper, we study the sampling of LLMs in three different settings. We observe that the sample has a shift from what is statistically likely towards an ideal value under some value system. This causes the LLM to show a value bias. We show that the LLM has an implicit value bias for concepts. We also show that value bias manifests even for new concepts learnt in context (zero-shot). We compare the value bias of LLMs to humans and show that the two might not always align. As a final remark, we would like to emphasize that we do not intend to contribute to “humanizing” AI/ML/LLMs in the way we use terminology or models. Our contribution is intended to draw parallels in behaviour and perform evaluations, as our findings will have an impact on downstream tasks.